OpenStudy (anonymous):

Write the equation for a circle with its center at (2,-5) and a radius of 6.

OpenStudy (anonymous):

(x-a)^2 + (y-b)^2 = r^2 is the Standard form of a circle. The center of the equation is (a,b). If the center is at (2,-5) than a point on the outside of the circle would be (2,1). These are your x and y values, respectively. So the equation would look like (2-2)^2 + (1-(-5))^2=6^2. Simplified, it is (2-2)^2 + (1+5)^2=36